# AICO: Feature Significance Tests for Supervised Learning

### v3

@@ B.3 Hybrid approach
Conditioning and grouping can be combined. For example, suppose we want to analyze a feature’s impact on model performance over a period of time. A long period avoids short-term noise but renders single-period conditioning insufficient. A shorter period ensures the relevance of data but may rule out grouping. This can be addressed by first conditioning on a time frame of desirable length and then grouping the conditional data by units corresponding to each time-unit pair. As with conditioning, this procedure can be repeated across different time windows, leading to a rolling-window analysis with smoothed single-period results.
Another notable use case of the hybrid method is the analysis of per-unit consistency of feature performance: i.e., for each unit, how consistent across time is a feature’s impact on model performance? A longer window reduces short-term fluctuations; when the multi-period score function is an aggregation of some single-period scores, a longer window also introduces another interesting aspect: per-unit consistency.
For example, suppose we choose the multi-period score function to simply be the average of a single-period function, such as negative mean squared error. Then, it can be shown that the unit-level feature effect $\Delta_{i}$ is essentially the average of unit-time-level feature effects $\Delta_{i,t}$ and, therefore, embeds some degree of consistency across time. The longer the time window, the higher the degree of consistency imposed, as the bar is set higher.

### v4

@@ B.3 Hybrid approach
Conditioning and grouping can be combined. For instance, suppose we wish to analyze a feature’s impact on model performance over an extended period. A longer period helps smooth short-term noise but makes single-period conditioning insufficient. Conversely, a shorter period preserves data relevance but may preclude grouping. This limitation can be addressed by first conditioning on a time frame of suitable length and then grouping the conditioned data by units corresponding to each time–unit pair. As with conditioning alone, this procedure can be repeated across overlapping time windows, yielding a rolling-window analysis with smoothed single-period results.
Another notable use case of the hybrid method is assessing per-unit consistency of feature performance—how consistent over time is a feature’s impact on model behavior for a given unit? A longer window mitigates short-term fluctuations, and when the multi-period score function aggregates single-period scores, it also captures an additional dimension of per-unit consistency.
For example, consider a multi-period score function defined as the average of a single-period function, such as the negative mean squared error. In this case, the unit-level feature effect $\Delta_{i}$ is effectively the average of the unit–time-level effects $\Delta_{i,t}$ and thus reflects a degree of consistency over time. The longer the time window, the stronger the imposed consistency, as a higher bar is set for variation in feature impact.